# Balanced Spanning Tree Distributions Have Separation Fairness

### v1

@@ 5.2 Technical Lemmas
Let $u$ be the blue vertex and let $R$ be the region that $u$ belongs to. If the vertex above $u$ is red (or does not exist), then $u$ is disposable from observation 19 and we are done. Similarly for the vertex below $u$. Thus, we assume that both these vertices (say, $v,w$) are blue. Let $R_{v}$, $R_{w}$ be the sets of vertices in $R\setminus\{u\}$ reachable from $v$ and $w$ respectively in $R\setminus\{u\}$. If $R_{v}=R_{w}=R\setminus\{u\}$, then we are done and $u$ is disposable.
@@
This follows from a simple case analysis. If the highlighted vertex is adjacent to a border, then it looks like Fig. 18a (where the unmarked black vertex may not exist). The blue vertex has degree at most 11 in the induced graph of blue vertices, so it is disposable from observation 19. Otherwise, the structure resembles Fig. 17b. If the 3 black vertices in Fig. 17b are all blue (as in Fig. 18b), then removing the vertex cannot affect the region. Similarly if exactly one neighbor of the highlighted vertex is blue (like Fig. 18c), then it is disposable from observation 19. If none of its neighbors are blue, then it forms a singleton region as in Fig. 18d and is hence, disposable.
@@ 6 Conclusion
Let $p_{k}$ denote the probability that the spanning tree distribution on the $m\times n$ gird samples a 2-partition where the smaller part has size $k$. Then for $i\in\{1,2,3\}$, $p_{mn/2}=\Theta(p_{mn/2+i})$.
@@ C.1 Cases 5 and 6
The highlighted vertex in Fig. 22c is red. In this case, the highlighted vertex in Fig. 23a must have a blue vertex above it. Otherwise, it is disposable (from observation 19) and we can flip it to red. But after this step, $u$ becomes disposable regardless of which case we started from, and $u$ can also be flipped to red to give a +2 partition. If this does not happen, then the picture must be as in Fig. 23b. Now, we flip the color of $u$ and the vertex above it, giving us the coloring in Fig. 23c. This creates two blue regions, one of which is an island, from Lemma 25. We also know that the blue vertices to the left of $u$ and 2 rows above $u$ are in different regions.
@@
If the highlighted vertex in Fig. 22c is blue, we have the coloring in Fig. 27a. If the vertex below $v$ was blue, then $v$ would be disposable from Observation 19, and we can flip it to blue. In this case, we get a +1 partition. Thus, the vertex below $v$ must be red. A similar argument can be made if the vertex to the right of $v$ is blue, or if the vertex to the right of $v$ is blue, but the highlighted vertex in Fig. 27b was red. In short, $v$ must be in 11Case 6. We now apply the elbow lemma (Lemma 24) as in Steps 2 and 3 (Figs. 22b and 22c) starting from the highlighted vertex in Fig. 27b since it is symmetric. If the neighborhood of $u$ was in 11Case 5, then we have the coloring in Fig. 27c. We note that this case is precisely the symmetric version of the previous subcase, and we can repeat the steps starting from Fig. 23a. The only remaining case is when the neighborhood of $u$ is in 11Case 6, just like the neighborhood of $v$. We now have the coloring in Fig. 28a. We have two cases based on the color of the highlighted vertex in Fig. 28a.
@@ D.5 Cases 10 and 11
If the two highlighted vertices in Fig. 68a are in different regions, then they form a 1-thin structure that can be flipped to produce a +0 partition. If they are in the same region, then we examine the three highlighted vertices in Fig. 68b. In the case where any of them are blue (from either region), we can resolve a 1-thin structure or 2-thin structure to form a +0 partition or a -1 partition respectively. Thus, we assume that the vertices are red. Next, we examine the two highlighted vertices in Fig. 68c. If the right vertex is red (or does not exist), then we can flip the vertex below $u$ to obtain a +2 partition. If the left vertex is blue, then we can flip it to red since it is disposable (observation 19). Thus, we assume that the left vertex is red and the right vertex is blue.

### v2

@@ 5.2 Technical Lemmas
Let $u$ be the blue vertex and let $R$ be the region that $u$ belongs to. If the vertex above $u$ is red (or does not exist), then $u$ is disposable from observation 22 and we are done. Similarly for the vertex below $u$. Thus, we assume that both these vertices (say, $v,w$) are blue. Let $R_{v}$, $R_{w}$ be the sets of vertices in $R\setminus\{u\}$ reachable from $v$ and $w$ respectively in $R\setminus\{u\}$. If $R_{v}=R_{w}=R\setminus\{u\}$, then we are done and $u$ is disposable.
@@
This follows from a simple case analysis. If the highlighted vertex is adjacent to a border, then it looks like Fig. 18a (where the unmarked black vertex may not exist). The blue vertex has degree at most 11 in the induced graph of blue vertices, so it is disposable from observation 22. Otherwise, the structure resembles Fig. 17b. If the 3 black vertices in Fig. 17b are all blue (as in Fig. 18b), then removing the vertex cannot affect the region. Similarly if exactly one neighbor of the highlighted vertex is blue (like Fig. 18c), then it is disposable from observation 22. If none of its neighbors are blue, then it forms a singleton region as in Fig. 18d and is hence, disposable.
@@ 6 Conclusion
Let $p_{j}$ denote the probability that the spanning tree distribution on the $m\times n$ gird samples a 2-partition where the smaller part has size $j$. Then for $i\in\{1,2,3\}$, $p_{mn/2}=\Theta(p_{mn/2-i})$.
@@
Acknowledgements
We thank Jamie Tucker-Foltz for helpful discussions and comments.
@@ C.1 Cases 5 and 6
The highlighted vertex in Fig. 22c is red. In this case, the highlighted vertex in Fig. 23a must have a blue vertex above it. Otherwise, it is disposable (from observation 22) and we can flip it to red. But after this step, $u$ becomes disposable regardless of which case we started from, and $u$ can also be flipped to red to give a +2 partition. If this does not happen, then the picture must be as in Fig. 23b. Now, we flip the color of $u$ and the vertex above it, giving us the coloring in Fig. 23c. This creates two blue regions, one of which is an island, from Lemma 28. We also know that the blue vertices to the left of $u$ and 2 rows above $u$ are in different regions.
@@
If the highlighted vertex in Fig. 22c is blue, we have the coloring in Fig. 27a. If the vertex below $v$ was blue, then $v$ would be disposable from Observation 22, and we can flip it to blue. In this case, we get a +1 partition. Thus, the vertex below $v$ must be red. A similar argument can be made if the vertex to the right of $v$ is blue, or if the vertex to the right of $v$ is blue, but the highlighted vertex in Fig. 27b was red. In short, $v$ must be in 11Case 6. We now apply the elbow lemma (Lemma 27) as in Steps 2 and 3 (Figs. 22b and 22c) starting from the highlighted vertex in Fig. 27b since it is symmetric. If the neighborhood of $u$ was in 11Case 5, then we have the coloring in Fig. 27c. We note that this case is precisely the symmetric version of the previous subcase, and we can repeat the steps starting from Fig. 23a. The only remaining case is when the neighborhood of $u$ is in 11Case 6, just like the neighborhood of $v$. We now have the coloring in Fig. 28a. We have two cases based on the color of the highlighted vertex in Fig. 28a.
@@ D.5 Cases 10 and 11
If the two highlighted vertices in Fig. 68a are in different regions, then they form a 1-thin structure that can be flipped to produce a +0 partition. If they are in the same region, then we examine the three highlighted vertices in Fig. 68b. In the case where any of them are blue (from either region), we can resolve a 1-thin structure or 2-thin structure to form a +0 partition or a -1 partition respectively. Thus, we assume that the vertices are red. Next, we examine the two highlighted vertices in Fig. 68c. If the right vertex is red (or does not exist), then we can flip the vertex below $u$ to obtain a +2 partition. If the left vertex is blue, then we can flip it to red since it is disposable (observation 22). Thus, we assume that the left vertex is red and the right vertex is blue.